OpenStudy (anonymous):

Essay: Show all work. The area of a rectangular swimming pool is given by 3x^2+11x+6 ft^2 . One side length of the pool is given by 3x + 2 feet. What is an algebraic expression for the other side length of the pool? Simplify this, and include correct units as part of your answer.

OpenStudy (hoblos):

3x^2+11x+6 =(3x + 2)(x+3) => the other side length is x+3 ft

OpenStudy (anonymous):

\[\Large \begin{array}{l} A = lw\\ l = 3x + 2\\ A = (3x + 2)w = 3{x^2} + 11x + 6\\ w = \frac{{3{x^2} + 11x + 6}}{{(3x + 2)}}\\ = \frac{{(3x + 2)(x + 3)}}{{(3x + 2)}}\\ = x + 3 \end{array}\]

OpenStudy (anonymous):

essay? lol

OpenStudy (anonymous):

Shortest essay I've ever written ;)

OpenStudy (anonymous):

how many words you have used?

OpenStudy (anonymous):

Do co-efficients count as words? :-P

OpenStudy (anonymous):

lol thank you for the help.

OpenStudy (anonymous):

fsm perfect